Problem 8-25 Two-Stage Dividend Growth Model [LO1]

Navel County Choppers Inc. is experiencing rapid growth. The company expects dividends to grow at 16 percent per year for the next 8 years before leveling off at 6 percent into perpetuity. The required return on the company’s stock is 15 percent. If the dividend per share just paid was $2.12, what is the stock price? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  Stock price $   
Answer #1

Stock Price is equal to the present value of all future dividends

Year Dividend Present value factor Present value
1 2.4592 0.869565217 2.138434783
2 2.852672 0.756143667 2.157029868
3 3.30909952 0.657516232 2.175786649
4 3.838555443 0.571753246 2.194706533
5 4.452724314 0.497176735 2.213790938
6 5.165160204 0.432327596 2.233041294
7 5.991585837 0.37593704 2.252459044
8 6.950239571 0.326901774 2.272045644
Terminal Value 81.85837717 0.326901774 26.7596487
Value 44.39694345

Hence, Stock Price = $44.4

Α 1 Year 2 1 3 2 4 3 5 4 65 7 6 8 7 98 10 Terminal Value 11 Dividend =2.12*1.16 =B2*1.16 =B3*1.16 =B4*1.16 =B5*1.16 =B6*1.16